# Continuous K-space Recovery Network with Image Guidance for Fast MRI Reconstruction

### v1

@@ Abstract
Magnetic resonance imaging (MRI) is a crucial tool for clinical diagnosis while facing the challenge of long scanning time. To reduce the acquisition time, fast MRI reconstruction aims to restore high-quality images from the undersampled k-space. Existing methods typically train deep learning models to map the undersampled data to artifact-free MRI images. However, these studies often overlook the unique properties of k-space and directly apply general networks designed for image processing to k-space recovery, leaving the precise learning of k-space largely underexplored. In this work, we propose a continuous k-space recovery network from a new perspective of implicit neural representation with image domain guidance, which boosts the performance of MRI reconstruction. Specifically, (1) an implicit neural representation based encoder-decoder structure is customized to continuously query unsampled k-values. (2) an image guidance module is designed to mine the semantic information from the low-quality MRI images to further guide the k-space recovery. (3) a multi-stage training strategy is proposed to recover dense k-space progressively. Extensive experiments conducted on CC359, fastMRI, and IXI datasets demonstrate the effectiveness of our method and its superiority over other competitors.
Index Terms:
Magnetic resonance imaging (MRI), image reconstruction, implicit neural representation, transformer.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Magnetic resonance imaging (MRI) is widely used for clinical diagnosis because of its advantages of no radiation, high resolution, and satisfactory soft-tissue contrast. However, the MRI acquisition process is inherently slow due to both physiological and hardware limitations, as well as the sequential nature of obtaining measurements in the frequency domain (k-space) [1, 2]. The extended scan time required to obtain fully sampled MRI images causes patient discomfort and reduces accessibility to the modality. To speed up this process, accelerated MRI aims to reconstruct MRI images from undersampled k-space [3, 4, 5]. However, the aliasing artifacts caused by such insufficient sampling often affect the clinical diagnosis. Therefore, it becomes a significant challenge how to reduce the amount of k-space acquisition while maintaining or even improving MRI image quality.
Essentially, the aliasing artifacts arise because undersampling destroys the k-space’s integrity. Many methods ignore the importance of k-space recovery and focus on reconstruction only from low-quality MRI images [6, 7, 8, 9, 10]. To this end, several works consider recovering k-space via an external network as shown in Fig.1 (a) [11, 12, 13, 14, 15, 16, 17, 18]. Despite some progress, they neglect the unique properties inherent to the k-space and directly adopt general CNNs which are not suitable for k-space recovery. CNNs have an inductive bias that their kernels are shared across spatial positions. However, in k-space, spatial positions represent frequency components of sine and cosine functions, and identical patterns at different positions can represent entirely different information[19, 20, 21, 22]. Previous works overlook this specific characteristic of k-space and adopt conventional image processing networks for processing k-space, resulting in inferior results. Hence, designing a customized network for undersampled k-space recovery is the key to reconstructing satisfactory MRI images.
@@ I Introduction
Intuitively, a high-quality MRI image is generated from a continuous k-space where each coordinate has a specific signal. Therefore, the ideal reconstruction of k-space should be modeled as a function where arbitrary coordinates can be mapped into a k-value. Such a concept of continuous representation is consistent with implicit neural representation (INR), which emerged as a new paradigm for image super-resolution [23, 24]. Specifically, INR uses a neural network to model images as continuous functions of spatial coordinates instead of discrete pixel grids. The network takes spatial coordinates as input and predicts corresponding pixel values. After training on low-resolution data, the model generates high-resolution images by querying dense coordinates, preserving fine details and enabling arbitrary resolution output [23, 25].
@@
Inspired by this, we propose to boost the performance of MRI reconstruction by recovering complete k-space via implicit neural representation. However, simply introducing INR to MRI reconstruction still faces several challenges. (1) INR originates from the processing of natural images, where rich semantic information is provided. Compared to it, the k-space spectrum exhibits a simple structure, making such continuous recovery challenging. Therefore, more information needed to be provided to guide k-space reconstruction. (2) The practice of accelerated scanning leads to excessively sparse sampled points, directly recovering a dense k-space from such sparsely sampled points inevitably introduces over-smoothing and distortion. A more gentle recovery paradigm should be designed to progressively reconstruct the dense k-space.
@@
Based on these, we propose an Image-domain Guided K-space Recovery Network, named IGKR-Net. As shown in Fig.1 (b), we customize a k-space network from a new perspective of INR with image guidance for MRI reconstruction.
@@
Specifically, (1) We propose an INR-based encoder-decoder structure, directly processing signals in the k-space domain. To detail, we design an encoder based on the standard transformer including multi-head self-attention (MSA) and feed-forward network (FFN), to encode the sampled k-values and their corresponding coordinates. Thus, a continuous feature representation of the k-space spectrum is learned. Subsequently, we adopt standard transformer decoder layers consisting of multi-head cross-attention (MCA), MSA, and FFN, to dynamically query the unsampled k-values using their coordinates and thus reconstruct the k-space spectrum with high quality. (2) We design an image domain guidance module to mine semantic information from the undersampled MRI images and guide the above k-space recovery. (3) We propose a tri-attention refinement module to further provide more details in the image domain. (4) To mitigate the issues of over-smoothing and distortion caused by directly recovering dense k-space from sparsely sampled points, we introduce a multi-stage training strategy to obtain finer reconstruction results step-by-step.
Overall, the main contributions of our work are as follows:
@@
• An INR based encoder-decoder structure is proposed to directly process signals in the k-space. Benefiting from this structure, our work can reconstruct the high-quality k-space.
@@
• An image domain guidance module is designed to mine the information from undersampled MRI images and guide the k-space recovery. Thanks to the design of this, we provide rich semantic information for the challenging k-space recovery.
@@
• A multi-stage training strategy is introduced to progressively reconstruct MRI images from low resolution to high resolution. Thanks to this paradigm, we mitigate concerns of over-smoothing and distortion and obtain precise results.
@@
• Extensive experiments are conducted on three public datasets, CC359, fastMRI, and IXI, demonstrating the superiority of our method over previous reconstruction approaches.
II RELATED WORK
II-A Deep Learning for Accelerated MRI
@@
Traditional MRI techniques involve the acquisition of dense k-space data, resulting in extended scan time. In recent years, a variety of deep neural networks have been explored for accelerated MRI. The pipeline for these methods usually consists of three steps, i.e., roughly recovering the complete k-space from undersampled data by employing a zero-filling strategy, then generating low-quality MRI images via inverse fast Fourier transform (IFFT), and finally learning the mapping relationship between such low-quality MRI images and their corresponding high-quality ones by adapting advanced CNN architectures, including UNet [26], deep residual network [27], generative adversarial network [28], and deep cascaded network [29].
@@
Besides, benefiting from the capability of learning global information, transformers [30, 31, 32, 33] have recently achieved satisfactory performance for accelerated MRI. Feng et al. [34, 35] pioneered the introduction of the transformer in the field of MRI reconstruction, integrating MRI reconstruction with super-resolution in a multitask fashion. Huang et al. [36] proposed a parallel imaging coupled swin transformer-based model. To sum up, the data-driven nature of deep learning allows these methods to learn the mapping between low-quality input and artifact-free images. However, such networks operate only in the image domain, making it challenging to ensure the consistency of the k-space.
@@
To this end, several attempts have been made to incorporate physics information in this line of work, including enforcing k-space consistency directly after image enhancement or adding k-space consistency as an additional cost function term during training. For example, Hyun et al. [11] enhanced k-space consistency by directly replacing the generated k-values with the corresponding original ones. Yang et al. [12] considered incorporating an additional k-space consistency term into the loss function. Though effective at reducing artifacts, they fundamentally only focus on image-domain restoration, with k-space information used for coarse correction or consistency supervision. The images relied upon by such methods are severely compromised, and this damage is irreversible [37, 38]. Indeed, the continuity of the k-space is crucial for the recovery of high-quality MRI images. However, these approaches neglect this and result in inferior results.
II-B K-space Recovery
@@
Recently, some studies began to recover k-space. Taejoon Eo et al. [39] proposed KIKINet, using a combination of 4 different CNNs to operate on k-space, image, k-space, and image sequentially. Osvald Nitski et al. [40] proposed CDF-Net, using two UNet to operate on the image and k-space domain respectively, and combining them into an end-to-end framework. Wang et al. [41] introduced a compressed sensing equivariant imaging prior framework that combines a data preparation strategy for generalizable MRI reconstruction. Zhao et al. [42] proposed SwinGAN, consisting of a k-space generator and an image-domain generator that utilizes swin transformer as the backbone. Such approaches aimed to preserve and leverage the k-space information, achieving more accurate results. However, they all treated the k-space spectrums as natural images, directly employing general networks for feature extraction and spectrum reconstruction.
@@
Besides, in k-space based parallel imaging reconstruction, some methods such as GRAPPA [43], SPIRiT [44] and AC-LORAKS [45] used linear convolution to interpolate undersampled k-space data. They assumed that the k-space data has a shifted autoregressive structure and learned the scan-specific autoregressive relationships required to recover missing sample scans from the fully sampled auto-calibration (ACS) data. However, at high acceleration, these methods produce inherent noise amplification and degrade image quality. Recently, the structure of the linear GRAPPA method has been transformed into a nonlinear deep learning method called RAKI [46]. RAKI performs scan-specific k-space interpolation using CNNs trained on the ACS data, and provides an improvement in both image quality and noise reduction over traditional PI methods that use linear k-space interpolation. However, the availability of training databases with different scans and their interpretability hinders its adaptability.
@@
Figure 2: The overall architecture of the proposed IGKR Net. Given the undersampled k-space, we first encode the coordinates and k-values of the sampled points, and the resulting features are added and then fed into LRIT. In LRIT, we use the LR coordinates to query and get the recovered results, i.e., $\widehat{{K}}_{1}$ and $\widehat{{I}}_{1}$. Next, we send $\widehat{{I}}_{1}$ into IDGM and get $\widehat{{K}}_{2}$ and $\widehat{{I}}_{2}$ with the guidance of the low-quality image $I_{s}$. Then, the k-values and coordinates of $\widehat{{K}}_{2}$ are encoded and fed into HRIT, where $\widehat{{K}}_{3}$ and $\widehat{{I}}_{3}$ are obtained by querying HR coordinates. Finally, we design TARM to refine $\widehat{{I}}_{3}$, yielding the final output $\widehat{{K}}_{4}$ and $\widehat{{I}}_{4}$.
II-C Implicit Neural Representation
@@
Recently, implicit neural representation (INR) has been proposed as a method of continuous representation for various tasks [47, 48, 49, 50, 51, 52]. INR employs neural networks, typically coordinate-based multilayer merceptrons (MLPs), to establish mapping relationships between coordinates and their corresponding signal values [53, 54].
@@
Because of its continuous and precise representation, INR has been applied to various tasks of image enhancement [23, 24]. For example, LIIF [23] proposed a new framework for arbitrary-scale super-resolution using INR. Li et al. [25] proposed a novel adaptive local image function (A-LIIF) to alleviate structural distortions and ringing artifacts around edges. Motivated by the feature interpolation in LIIF, Wu et al. [55] designed an INR-based network for achieving 3D MRI super-resolution at arbitrary scales. In summary, 2D INR models an image as a continuous function using latent space interpolation, enabling precise querying of signal intensity for any given coordinates. This concept provides a novel approach for achieving high-quality recovery in k-space.
III METHODOLOGY
III-A Preliminary
III-A1 MRI Reconstrucion
@@
Let $K$ represent the complex-valued, fully sampled k-space acquired from the MRI scanner, accelerated MRI usually employs undersampling to acquire a reduced set of k-space, i.e., $K_{s}$, while a substantial portion of the k-space, i.e., $K_{us}$, is unsampled. Here, we simulate this undersampling process by the element-wise multiplication ($\otimes$) of $K$ with a two-dimensional mask $M$:
$K_{s}=M\otimes K,K_{us}=(1-M)\otimes K$ (1)
@@
Correspondingly, low-quality MRI image $I_{s}$ can be obtained, i.e., $I_{s}=IFFT(K_{s})$. To recover high-quality MRI image $\widehat{{I}}$, deep learning methods typically leverage extensive training data to establish a mapping relationship between $\widehat{{I}}$ and the sampled data $(I_{s},K_{s})$, which can be formulated as:
$\widehat{{I}}=f_{\theta}(I_{s},K_{s})$ (2)
@@
Here, $\theta$ is the parameters set of the deep neural network. In contrast to these methods, our IGKR-Net focuses on recovering $K_{us}$, which is the fundamental cause of MRI image blurring and is crucial for high-quality MRI reconstruction.

### v2

@@ Abstract
Magnetic resonance imaging (MRI) is a crucial tool for clinical diagnosis while facing the challenge of long scanning time. To reduce the acquisition time, fast MRI reconstruction aims to restore high-quality MRI images from the undersampled k-space. Existing methods typically train deep learning models to map the undersampled data to artifact-free MRI images. However, these studies often overlook the unique properties of k-space and directly apply general networks designed for image processing to k-space recovery, leaving the precise learning of k-space largely underexplored. In this work, we propose a continuous k-space recovery network from a new perspective of implicit neural representation with image domain guidance, which boosts the performance of MRI reconstruction. Specifically, (1) an implicit neural representation based encoder-decoder structure is customized to continuously query unsampled k-values. (2) an image guidance module is designed to mine the image-domain information from the low-quality MRI images to further guide the k-space recovery. (3) a multi-stage training strategy is proposed to recover dense k-space progressively. Extensive experiments conducted on CC359, fastMRI, IXI, and SKM-TEA datasets demonstrate the effectiveness of our method and its superiority over other competitors. Implementation code will be available in https://github.com/XiaoMengLiLiLi/IGKR.
keywords:
[…]
1 Introduction
Magnetic resonance imaging (MRI) is widely used for clinical diagnosis because of its advantages of no radiation, high resolution, and satisfactory soft-tissue contrast. However, the acquisition of dense k-space for accurate MRI images inherently leads to long time due to both physiological and hardware limitations (Scherzinger and Hendee, 1985; Doneva, 2020). The extended scan time causes patient discomfort and reduces accessibility to MRI. To speed up this process, accelerated MRI aims to reconstruct MRI images from undersampled k-space (Lustig et al., 2007; Gamper et al., 2008; Pruessmann et al., 1999). However, the aliasing artifacts caused by such insufficient sampling often affect the clinical diagnosis. Therefore, it becomes a significant challenge how to reduce the amount of k-space acquisition while maintaining or even improving MRI image quality.
@@ 1 Introduction
In recent years, deep learning has emerged as a powerful tool to tackle this problem. These methods usually learn the mapping between low-quality MRI images and their corresponding high-quality ones by adapting advanced neural network architectures (Guo et al., 2023; Meng et al., 2024; Hammernik et al., 2018; Aghabiglou et al., 2024; Velayudham et al., 2025). In other words, the main purpose of this methodology is to achieve MRI reconstruction purely in the image domain. Although they have succeeded in removing artifacts in the image domain, k-space recovery cannot be achieved, resulting in unsatisfactory reconstruction results.
Essentially, the aliasing artifacts arise because the accelerated undersampling destroys the k-space’s integrity. To this end, several works consider recovering k-space via an external network as shown in Figure.1 (a) (Eo et al., 2018; Zhou and Zhou, 2020; Ran et al., 2020; Zhao et al., 2023). Despite some progress, they neglect the unique properties inherent to the k-space and directly adopt general CNNs which are not suitable for k-space recovery. CNNs have an inductive bias that their kernels are shared across spatial positions. However, in k-space, spatial positions represent frequency components of sine and cosine functions, and identical patterns at different positions can represent entirely different information(Minhas and Oliver, 2022; Moore and Chung, 2017). Previous works overlook this specific characteristic and adopt general image networks for k-space processing, resulting in inferior results. Hence, designing a customized network for undersampled k-space recovery is the key to reconstructing satisfactory MRI images.
@@
Intuitively, a high-quality MRI image is generated from a continuous k-space where each coordinate has a specific signal. Therefore, the ideal reconstruction of k-space should be modeled as a function where arbitrary coordinates can be mapped into a k-value. Such a concept of continuous representation is consistent with implicit neural representation (INR), which emerged as a new paradigm for image super-resolution (Chen et al., 2021; Xu et al., 2021). INR uses a neural network to model images as continuous functions of spatial coordinates instead of discrete pixel grids. Specifically, the neural network takes low-resolution coordinates as input and predicts corresponding pixel values, thereby constructing a latent space. After that, the model generates high-resolution images by querying dense coordinates from the latent space, preserving fine details and enabling arbitrary resolution output (Chen et al., 2021; Li et al., 2022).
@@
Inspired by this, we propose to boost the performance of MRI reconstruction by recovering complete k-space via INR. However, simply introducing INR to MRI reconstruction still faces several challenges. (1) INR originates from the processing of natural images, where rich structure information is provided for effective latent space modeling. Compared to it, the k-space spectrum exhibits a simple structure, making such continuous recovery challenging. Therefore, more information needed to be provided to guide k-space reconstruction. (2) The practice of accelerated scanning leads to excessively sparse sampled points, directly recovering a dense k-space from such sparsely sampled points inevitably introduces over-smoothing and distortion. A more gentle recovery paradigm should be designed to progressively reconstruct the dense k-space.
@@
Based on these, we propose an Image-domain Guided K-space Recovery Network, named IGKR-Net. As shown in Figure.1 (b), we customize a k-space net from a new perspective of INR with image guidance for MRI reconstruction.
@@
Specifically, (1) We propose an INR-based encoder-decoder structure, directly processing signals in the k-space domain. To detail, we design an encoder based on the standard transformer including multi-head self-attention (MSA) and feed-forward network (FFN), to encode the sampled k-values and their corresponding coordinates. Thus, a continuous latent space of the k-space spectrum is learned. Subsequently, we adopt standard transformer decoder layers consisting of multi-head cross-attention (MCA), MSA, and FFN, to dynamically query the unsampled k-values using their coordinates and thus reconstruct the k-space spectrum with high quality. (2) We design an image domain guidance module to mine image-domain information from the undersampled MRI images and guide the above k-space recovery. (3) We propose a tri-attention refinement module to further provide more details in the image domain. (4) To mitigate the issues of over-smoothing and distortion caused by directly recovering dense k-space from sparsely sampled points, we introduce a multi-stage training strategy to obtain finer reconstruction results step-by-step.
The main contributions of our work are as follows:
@@
  • An implicit neural representation based encoder-decoder structure is proposed to directly process signals in the k-space. Benefiting from this structure, our work can reconstruct the high-quality k-space.
@@
  • An image domain guidance module is designed to mine information from undersampled MRI images. Due to the design of this, we provide rich information for the challenging k-space recovery.
@@
  • A multi-stage training strategy is introduced to progressively reconstruct MRI images from low resolution to high resolution. Benefiting from this, we mitigate concerns of over-smoothing and distortion and obtain precise results.
@@
  • Extensive experiments are conducted on both single-coil and multi-coil datasets, demonstrating the superiority of our method over previous approaches.
2 Related Work
2.1 Deep Learning for Accelerated MRI
@@
Traditional MRI involves the acquisition of dense k-space, resulting in long scan time. In recent years, many deep neural networks have been explored for accelerated MRI. The pipeline for these methods usually consists of three steps, i.e., roughly recovering the complete k-space from undersampled data via zero-filling, generating low-quality MRI images via inverse fast Fourier transform (IFFT), and learning the mapping relationship between such low-quality MRI images and their corresponding high-quality ones by adapting advanced CNN architectures, including UNet (Zbontar et al., 2018), deep residual network (Lee et al., 2018), generative adversarial network (Mardani et al., 2019), and deep cascaded network (Jo et al., 2017).
@@
Besides, benefiting from the capability of learning global information, transformers (Yang et al., 2024a, c, b) recently achieved satisfactory performance for accelerated MRI. Feng et al. (Feng et al., 2021, 2022) pioneered the introduction of the transformer in the field of MRI reconstruction, integrating MRI reconstruction with super-resolution in a multitask fashion. Huang et al. (Huang et al., 2022) proposed a parallel imaging coupled swin transformer-based model. Reconformer  (Guo et al., 2023) introduced a locally pyramidal but globally columnar architecture to perceive multi-scale representation at any stage for MRI image reconstruction. To sum up, the data-driven nature of deep learning allows these methods to learn the mapping between low-quality input and artifact-free images. However, such networks operate only in the image domain, making it challenging to ensure the consistency of the k-space.
@@
To this end, several attempts have been made to incorporate physics information in this line of work, including enforcing k-space consistency directly after image enhancement or adding k-space consistency as an additional cost function term during training. For example, Hyun et al. (Hyun et al., 2018) enhanced k-space consistency by directly replacing the generated k-values with the corresponding original ones. Yang et al. (Yang et al., 2017) considered incorporating an additional k-space consistency term into the loss function. Though effective at reducing artifacts, they fundamentally only focus on image-domain restoration, with k-space information used for coarse correction or consistency supervision. The images relied upon by such methods are severely compromised, and this damage is irreversible (Vassiliou et al., 2018; Ravishankar and Bresler, 2011). Indeed, the continuity of the k-space is crucial for the recovery of high-quality MRI images. However, these approaches neglected this and resulted in inferior results.
2.2 K-space Recovery
@@
Recently, some studies began to recover k-space. Taejoon Eo et al. (Eo et al., 2018) proposed KIKINet, using a combination of 4 different CNNs to operate on k-space, image, k-space, and image sequentially. Osvald Nitski et al. (Nitski et al., 2020) proposed CDF-Net, using two UNet to operate on the image and k-space domain respectively, and combining them into an end-to-end framework. Wang et al. (Wang et al., 2024) introduced a compressed sensing equivariant imaging prior framework that combines a data preparation strategy for generalizable MRI reconstruction. Zhao et al. Zhao et al. (2023) proposed SwinGAN, consisting of a k-space generator and an image-domain generator that utilizes swin transformer as the backbone. These approaches aimed to preserve and leverage the k-space information, achieving accurate results. However, they treated the k-space spectrums as natural images, directly employing general networks for feature extraction and spectrum reconstruction.
@@
Besides, as for k-space based parallel reconstruction, GRAPPA (Griswold et al., 2002), SPIRiT (Lustig and Pauly, 2010) and AC-LORAKS (Haldar, ) used linear convolution to interpolate undersampled k-space. They assumed that the k-space data has a shifted autoregressive structure and learned the scan-specific autoregressive relationships to recover missing sample scans from the fully sampled auto-calibration (ACS) data. However, at high acceleration, they produce inherent noise amplification. Recently, the structure of the linear GRAPPA method has been transformed into a nonlinear deep learning method called RAKI (Akçakaya et al., 2019). However, the availability of training ACS databases with different scans hinders its adaptability.
@@
Figure 2: The architecture of the our IGKR-Net. Given the undersampled k-space, we first encode the coordinates $C_{s}$ and k-values $K_{s}$ of the sampled points, and the resulting features are added and then fed into LRIT. In LRIT, we use the LR coordinates to query and get the recovered results, i.e., $\widehat{{K}}_{1}$ and $\widehat{{I}}_{1}$. Next, we send $\widehat{{I}}_{1}$ into IDGM and get $\widehat{{K}}_{2}$ and $\widehat{{I}}_{2}$ with the guidance of the low-quality image $I_{s}$. Then, the k-values $K_{s}^{{}^{\prime}}$ and coordinates $C_{s}^{{}^{\prime}}$ of $\widehat{{K}}_{2}$ are encoded and fed into HRIT, where $\widehat{{K}}_{3}$ and $\widehat{{I}}_{3}$ are obtained by querying HR coordinates. Finally, we design TARM to refine $\widehat{{I}}_{3}$, yielding the final output $\widehat{{K}}_{4}$ and $\widehat{{I}}_{4}$.
Figure 3: Our transformer based Encoder (a) and Decoder (b).
2.3 Implicit Neural Representation
@@
Recently, implicit neural representation (INR) has been proposed as a method of continuous representation for various tasks (Sitzmann et al., 2020; Grattarola and Vandergheynst, 2022; Chen et al., 2022; Chen and Zhang, 2019; Mildenhall et al., 2021; Rahaman et al., 2019). INR employs neural networks, typically coordinate-based multilayer merceptrons (MLPs), to establish mapping relationships between coordinates and their corresponding signal values (Zheng et al., 2024; Chen et al., 2024; Li et al., 2025).
@@
Because of its continuous and precise representation, INR has been applied to various tasks of image enhancement (Chen et al., 2021; Xu et al., 2021). For example, LIIF (Chen et al., 2021) proposed a new framework for arbitrary-scale super-resolution using INR. Li et al. (Li et al., 2022) proposed a novel adaptive local image function (A-LIIF) to alleviate structural distortions and ringing artifacts around edges. Motivated by the feature interpolation in LIIF, Wu et al. (Wu et al., 2022) designed an INR-based network for 3D MRI super-resolution at arbitrary scales. In summary, 2D INR models an image as a continuous function using latent space, enabling precise querying of signal intensity for any given coordinates. This concept provides a novel approach for achieving high-quality recovery in k-space.
3 Methodology
3.1 Preliminary
3.1.1 MRI Reconstrucion
@@
Let $K$ represent the complex-valued, fully sampled k-space acquired from the MRI scanner, accelerated MRI usually employs undersampling to acquire a reduced set of k-space, i.e., $K_{s}$, while a substantial portion of the k-space, i.e., $K_{us}$, is unsampled. Here, we simulate this undersampling process by the element-wise multiplication ($\otimes$) of $K$ with a two-dimensional mask $M$ as follows:
$K_{s}=M\otimes K,K_{us}=(1-M)\otimes K.$ (1)
@@
Low-quality MRI image $I_{s}$ can be obtained, i.e., $I_{s}=IFFT(K_{s})$. To recover high-quality MRI image $\widehat{{I}}$, deep learning methods typically leverage extensive training data to establish a mapping relationship between $\widehat{{I}}$ and the sampled data $(I_{s},K_{s})$, which can be formulated as follows:
$\widehat{{I}}=f_{\theta}(I_{s},K_{s}),$ (2)
@@
where $\theta$ is the parameters set of the deep neural network. Different from these methods, our IGKR-Net focuses on recovering $K_{us}$, which is the fundamental cause of MRI image blurring and is crucial for high-quality MRI reconstruction.
3.1.2 Implicit Neural Representation
In contrast to the typical representation of images using discrete pixel arrays, implicit neural representation (INR) models the mapping relationship between image coordinates $C$ and their corresponding intensity values $i$ as a function, which can be formulated as follows:
$i=f_{\theta}(C,V_{c}),$ (3)
where $V_{c}$ represents the feature vector specific to the pixel at spatial coordinates $C$. $V_{c}$ is obtained by establishing a latent space by training an encoder. Function $f_{\theta}$ is fitted using a neural network and is continuous to the spatial coordinate system. Therefore, once $f_{\theta}$ is well-fitted, intensity values can be queried at arbitrary coordinates, thereby achieving a continuous visual representation of the image.
3.1.3 Continuous K-space Recovery Network via INR
As described in equation (1), it is common to undersample the k-space for accelerated MRI. The correct recovery of the unsampled k-values, i.e., $K_{us}$, is crucial for high-quality MRI reconstruction. Inspired by the continuous image modeling of INR, we customize a k-space recovery net. Specifically, $K_{us}$ can be queried via their coordinates $C_{us}$:
$K_{{us}}=f_{\theta}(C_{{us}},V_{C_{us}}),$ (4)
$K_{{us}}=\Phi_{Dec}(C_{{us}},\Phi_{Enc}(K_{s},C_{s})),$ (5)
where the encoder $\Phi_{Enc}$ and decoder $\Phi_{Dec}$ are implemented through the transformer attention mechanism.
3.2 Network Architecture
As shown in Figure.2, the overall architecture of our IGKR-Net can be divided into two parts, i.e., the primary part directly processing in k-space domain (denoted as red dash lines), and the auxiliary part operating on image domain (denoted as blue dash lines). The k-space part comprises LRIT and HRIT, to reconstruct the k-space from low resolution (LR) to high resolution (HR) progressively, while the image part comprises IDGM and TRAM for image guidance and final refinement, respectively.
Specifically, given undersampled k-space $K\in R^{2\times H\times W}$ as input, we obtain the set of k-values $K_{s}$ and the corresponding coordinates $C_{s}$, from the sampled points. Then, $K_{s}$ and $C_{s}$ are encoded separately, and the resulting features are summed to form a vector sequence $\mathcal{V}$, which is used as the input of LRIT. LRIT first uses a transformer-based encoder to learn a latent space $\mathcal{S}$, and then recover a complete while LR k-space $\widehat{{K}}_{1}\in R^{2\times\frac{H}{2}\times\frac{W}{2}}$ by querying pre-designed LR coordinates through the decoder. Subsequently, we get $\widehat{{I}}_{1}$ through IFFT and send it into IDGM to deeply mine information from the corrupted image $I_{s}$ and guide the k-space recovery. We fuse $I_{s}$ with $\widehat{{I}}_{1}$, obtaining $\widehat{{I}}_{2}\in R^{2\times H\times W}$ and $\widehat{{K}}_{2}\in R^{2\times H\times W}$. Next, HRIT uses the recovered dense $\widehat{{K}}_{2}$ to further establish a continuous latent space $\mathcal{S}^{{}^{\prime}}$, recovering complete and HR $\widehat{K}_{3}$ and its corresponding $\widehat{{I}}_{3}$ using the HR coordinates. Finally, we design TARM to refine $\widehat{{I}}_{3}$ in the image domain, resulting in the final output, i.e., $\widehat{{I}}_{4}$ and $\widehat{{K}}_{4}$.